obtain the  ΔH° of the following:


  1. Zn(s) + 2H+(aq) → Zn2+(aq) + H2(g)
1
Expert's answer
2021-09-04T02:33:13-0400

ΔH=ΔE+PΔV

or ΔE=ΔH−PΔV

PΔV=1 atm×22.4 L=22.4 L atm

=22.4×101.3 J=2307 J=2.31 kJ

∴ΔE=−154.4−2.31=−156.71 kJ